OpenStudy (anonymous):

A 2,500 kg car moves at 15 m/s2; 15 m/s2 is the ___________ of the car

OpenStudy (anonymous):

Acceleration. Remember, units are ESSENTIAL to definition. \[\frac{ dx }{ dt } = velocity = \frac{ m }{ s }\] \[\frac{ dv }{ dt } = acceleration = \frac{ m }{ s^2}\]